# Improved GUI Grounding via Iterative Narrowing

### v5

@@ 1 Introduction
GUI Grounding is the task of identifying the visual location on an interface image given a natural language query [4, 10]. This task is essential for improving the capabilities of Vision Language Model (VLM) agents, enabling them to better understand and interact with graphical user interfaces [11]. While general VLMs, such as GPT-4V, have demonstrated high performance across a variety of visual-linguistic tasks, their accuracy in GUI grounding tasks remains suboptimal [4, 11].
Existing solutions have attempted to bridge this performance gap by fine-tuning VLMs for GUI grounding, resulting in improvements over baseline models [5, 4, 10, 6]. However, these approaches often come with the need for extensive data and re-training.
@@
We evaluate the performance of our method on the ScreenSpot [4] benchmark. Our findings demonstrate that this iterative approach leads to substantial performance gains, providing a straightforward yet effective solution for enhancing GUI grounding capabilities.
@@ 2.1 Grounding Models
Cheng et al. [4] trained SeeClick, a finetuned grounding model, and introduced the ScreenSpot benchmark, demonstrating the benefits of grounding pre-training. Gou et al. [5] achieved up to 20% accuracy gains with a pixel-level model trained on 10 million GUI elements. Wu et al. [10] developed OS-Atlas models using a 13-million GUI element dataset, improving performance across benchmarks. Lin et al. [7] proposed ShowUI, a lightweight vision-language-action model optimized for GUI tasks using innovative token selection and streaming methods.
@@ 2.2 GUI Agents
Xie et al. [11] introduced the OSWorld benchmark to evaluate GUI task agents, emphasizing grounding in screenshot-only scenarios. Wu et al. [10] enhanced task performance by integrating OS-Atlas-Base-7B with GPT-4o. Hong et al. [6] addressed GUI grounding with CogAgent, designed as both a grounding model and visual agent.
2.3 Iterative Reasoning and Localization
Chen et al. [1] introduced a visual reasoning framework combining spatial memory and graph reasoning modules to iteratively refine predictions through spatial and semantic interactions. Similarly, Sun et al. [8] proposed an iterative shrinking approach for referring expression grounding using reinforcement learning, which parallels our method’s reliance on VLM predictions for target refinement.
@@ 4.1 Setup
We evaluated our method on the ScreenSpot [4] benchmark, a comprehensive assessment tool for single-step GUI grounding across multiple platforms. The ScreenSpot benchmark is structured into three primary categories: mobile, web, and desktop, each designed to reflect common user interface environments. Furthermore, each category is divided into two subcategories based on the type of target element: text or icon.
@@ 4.3 Weaknesses & Observations
Figure 2: An example of a context-dependent failure case. At iteration $k<n$ (top image), the cropping leads to the loss of crucial contextual information—specifically, the association between each star button and its corresponding project. Thus, resulting in the wrong UI element being selected (bottom image).
Our method improves precision but at the cost of underperforming in context-dependent scenarios. This limitation likely explains the greater gains observed in generalist VLMs (e.g., InternVL-2-4B [3, 2], Qwen2-VL-7B [9]) versus OS-Atlas [10] and ShowUI [7], which already has strong precision abilities from training. Addressing this limitation by maintaining context throughout the iterations could further enhance performance for all models.
4.4 Preliminary Solutions and Future Work
To address the loss of contextual information, we conducted preliminary experiments where the model was provided with both the entire screenshot and the current crop during each iteration. This approach was inspired by the work of Chen et al. [1], which introduced a mechanism to leverage both local and global modules that iteratively exchange information for enhanced reasoning. Our goal was to enable the model to retain a global understanding while refining its focus through local crops.
However, during these initial experiments, we observed that the Visual Language Model (VLM) frequently confused the local crop with the global context image. This confusion led to incorrect coordinate predictions, as the model sometimes generated coordinates referencing the entire global image rather than accurately focusing on the local context intended for refinement.
@@ 5 Conclusion
In this work, we introduced a visual prompting framework designed to perform an iterative narrowing mechanism. As a result, this enhances the GUI grounding capabilities of Vision-Language Models (VLMs). Through iterative refinement of predictions, our method improves accuracy by narrowing the focus on progressively smaller regions, allowing models to more precisely identify visual elements on a graphical user interface. Evaluation on the ScreenSpot [4] benchmark demonstrated that our method improves performance, especially for generalist VLMs like InternVL-2-4B [2, 3] and Qwen2-VL-7B [9], where substantial improvements were observed compared to baselines. However, the method shows limitations when handling spatially distant contextual cues, which affects performance in context-dependent scenarios.

### v6

@@ 1 Introduction
GUI Grounding is the task of identifying the visual location on an interface image given a natural language query Cheng et al. (2024); Wu et al. (2024). This task is essential for improving the capabilities of Vision Language Model (VLM) agents, enabling them to better understand and interact with graphical user interfaces Xie et al. (2024). While general VLMs, such as GPT-4V, have demonstrated high performance across a variety of visual-linguistic tasks, their accuracy in GUI grounding tasks remains suboptimal (Cheng et al., 2024; Xie et al., 2024).
Existing solutions have attempted to bridge this performance gap by fine-tuning VLMs for GUI grounding, resulting in improvements over baseline models (Gou et al., 2024; Cheng et al., 2024; Wu et al., 2024; Hong et al., 2023). However, these approaches often come with the need for extensive data and re-training.
@@
We evaluate the performance of our method on the ScreenSpot (Cheng et al., 2024) benchmark. Our findings demonstrate that this iterative approach leads to substantial performance gains, providing a straightforward yet effective solution for enhancing GUI grounding capabilities.
@@ 2 Related Work
2.1 Iterative Reasoning and Localization
Iterative refinement strategies have been explored in multiple contexts, aiming to progressively narrow down the target location. Wu et al. Wu and Xie (2023) proposed SEAL, an LLM-guided visual search framework designed for multimodal large language models in high-resolution natural image understanding, particularly for visual question answering tasks. SEAL integrates a dedicated visual search model, specialized localization decoders, and a Visual Working Memory to actively recover missing visual details, requiring additional model inference beyond the base VLM. In contrast, our iterative narrowing method is developed specifically for precise GUI element grounding, is training-free, and reuses the same VLM across successive refinement steps. Chen et al. Chen et al. (2018) introduced a visual reasoning framework that combines spatial memory with graph reasoning to iteratively refine predictions through spatial and semantic cues. Similarly, Sun et al. Sun et al. (2021) proposed an iterative shrinking approach for referring expression grounding and relies on reinforcement learning to determine the shrinking strategy, whereas we perform refinement purely through repeated model predictions without additional training or policy learning.
@@ 2.2 Grounding Models
Recent work on GUI grounding has explored both large-scale pretraining and task-specific model design. Cheng et al. (Cheng et al., 2024) trained SeeClick, a finetuned grounding model, and introduced the ScreenSpot benchmark, highlighting the benefits of grounding-specific pretraining. Gou et al. (Gou et al., 2024) achieved up to 20% accuracy gains using a pixel-level model trained on 10 million GUI elements. Wu et al. (Wu et al., 2024) developed OS-Atlas models with a 13-million GUI element dataset, delivering consistent improvements across multiple benchmarks. Lin et al. (Lin et al., 2024) proposed ShowUI, a lightweight vision-language-action model optimized for GUI tasks through efficient token selection and streaming methods.
@@ 2.3 GUI Agents
Beyond grounding in static scenarios, other works have explored GUI agents that perform end-to-end interaction. Xie et al. Xie et al. (2024) introduced the OSWorld benchmark to evaluate such agents, emphasizing grounding in screenshot-only settings. Wu et al. Wu et al. (2024) improved task performance by combining OS-Atlas-Base-7B with GPT-4o  OpenAI (2024) for enhanced reasoning and perception. Hong et al. Hong et al. (2023) presented CogAgent, which serves as both a grounding model and an interactive visual agent.
@@ 4.1 Setup
We evaluated our method on the ScreenSpot (Cheng et al., 2024) benchmark, a comprehensive assessment tool for single-step GUI grounding across multiple platforms. The ScreenSpot benchmark is structured into three primary categories: mobile, web, and desktop, each designed to reflect common user interface environments. Furthermore, each category is divided into two subcategories based on the type of target element: text or icon.
@@ 4.3 Weaknesses & Observations
Figure 2: An example of a context-dependent failure case. At iteration $k<n$ (top image), the cropping leads to the loss of crucial contextual information. Specifically, the association between each star button and its corresponding project. Thus, resulting in the wrong UI element being selected (bottom image).
Our method improves precision but at the cost of underperforming in context-dependent scenarios. This limitation likely explains the greater gains observed in generalist VLMs (e.g., InternVL-2-4B Chen et al. (2024, 2023), Qwen2-VL-7B Wang et al. (2024)) versus OS-Atlas Wu et al. (2024) and ShowUI Lin et al. (2024), which already has strong precision abilities from training. Addressing this limitation by maintaining context throughout the iterations could further enhance performance for all models.
4.4 Potential Solutions and Future Work
To address the loss of contextual information, we conducted preliminary experiments where the model was provided with both the entire screenshot and the current crop during each iteration. The goal was to enable the model to retain a global understanding while refining its focus through local crops.
However, during these initial experiments, we observed that the VLM frequently confused the local crop with the global context image. This confusion led to incorrect coordinate predictions, as the model sometimes generated coordinates referencing the entire global image rather than accurately focusing on the local context intended for refinement.
@@ 5 Conclusion
In this work, we introduced a visual prompting framework designed to perform an iterative narrowing mechanism. As a result, this enhances the GUI grounding capabilities of Vision-Language Models (VLMs). Through iterative refinement of predictions, our method improves accuracy by narrowing the focus on progressively smaller regions, allowing models to more precisely identify visual elements on a graphical user interface. Evaluation on the ScreenSpot Cheng et al. (2024) benchmark demonstrated that our method improves performance, especially for generalist VLMs like InternVL-2-4B Chen et al. (2023, 2024) and Qwen2-VL-7B Wang et al. (2024), where substantial improvements were observed compared to baselines. However, the method shows limitations when handling spatially distant contextual cues, which affects performance in context-dependent scenarios.